OpenStudy (anonymous):

Create your own unique quadratic equation in the form y = ax^2 + bx + c. Use complete sentences and show all work to determine the following: Does the graph open up or down? How do you know? Explain whether the graph has a maximum or minimum point. Find the vertex and x-intercepts of the graph.

OpenStudy (anonymous):

Form: \(y=ax^2+bx+c\). For \(a>0\), this opens up, since a positive coefficient in front of the leading term means the limits go to \(+\infty\). Imagine a parabola in your head. If the graph opens up, the maximum is undefined at \(\infty\), while the minimum is a specific point, right? The last part should be relatively straightforward. For ease, I suggest you pick \((x+1)^2\). And laugh as you skip most of the work.